OpenStudy (anonymous):

convert the mixed number 5 1/3

OpenStudy (ryan123345):

You multiply the denominator times the whole number and add that to the numerator.

OpenStudy (ryan123345):

\[3 \times 5 + 1\]

OpenStudy (ryan123345):

The answer to that will give you your numerator and you just keep the same denominator which is 3 ....

OpenStudy (anonymous):

16/3

OpenStudy (ryan123345):

Good Job! :)

OpenStudy (anonymous):

but how do this reduce to 1 2/3 is what im not understanding

OpenStudy (ryan123345):

You are not understanding how to put that into improper?

OpenStudy (anonymous):

yes

OpenStudy (ryan123345):

Do you have any guesses what the answer would be?

OpenStudy (anonymous):

this is the problem 12in=1ft+4ft=5ft*1yd/3ft=1 2/3 I do not understand they reduced this problem to 1 2/3

OpenStudy (unklerhaukus):

to convert 5 into thirds, multiply it by 3/3 so 5 = 5*3/3 = 15/3 this means 5+1/3 =15/3+1/3 ..

OpenStudy (unklerhaukus):

=16/3 as you said.

OpenStudy (unklerhaukus):

for 1 2/3 again we want thirds so change 1 by multiplying by 3/3 1*3/3 = 3/3 so 3/3 + 2/3 = (3+2)/3 =